Perform the calculation. 69°20 - 35°48 o 69°20 - 35°48 =
Answer #1

Given Now convert 69° 20 35°48 minute in degree 69°20 I minute - 1 1 degree? 699 20 60 35°48 11 = 690 + 7 = (63+2) sº +40